The wheel, rotating at uniformly accelerated speed, after a time t =1min after the start of rotation, acquires a frequency on n=720rmp. Find the angular acceleration e of the wheel and number of revolutions N of the wheel during this time
1
Expert's answer
2019-04-23T11:12:19-0400

The angular velocity

"\omega=\omega_0+\varepsilon t"

So, frequency

"2\pi n=2\pi n_0+\varepsilon t"

Finally, the angular acceleration

"\varepsilon=\frac{2\pi n_0}{t}=\frac{2\pi \times\frac{720\:\rm{rev}}{60\:\rm{s}}}{60\:\rm{s}}=0.4\pi\:\rm{s^{-2}}=1.256\:\rm{s^{-2}}"

The number of revolutions

"N=\frac{\frac{\varepsilon t^2}{2}}{2\pi}=\frac{0.4\times 60^2}{4}=360\:\rm{rev}"